# Macroscopic properties of equity markets: stylized facts and portfolio performance

### v1

@@ 1. Introduction
In this paper we study empirically a collection of macroscopic properties of equity markets which are highly relevant in the management of large equity portfolios but are not adequately addressed by the aforementioned literature. In a nutshell, a macroscopic quantity or property of an equity market is one which depends on all or a majority of the stocks in the market. The most basic example is the total market capitalization, or overall performance, of the market which is well represented by a capitalization-weighted market index such as the S&P 500. The index often serves as a benchmark for equity portfolios, both passive and active, and may be approximated by tradeable assets such as exchange-traded funds. Our terminology is motivated by statistical physics in which one considers macroscopic quantities such as volume, temperature and pressure. To take our crude physical analogy a bit further, suppose we think of an equity market as a galaxy and each stock as a star. Then the market index is a proxy of the total mass, but there are other interesting macroscopic quantities, such as the shape, spiral and rotation of the galaxy, which affect and are affected by the stars, and are much less understood. [^1] There are also other galaxies (markets) but they are beyond the scope of this paper. From the viewpoint of, say, the manager of an equity portfolio which aims to outperform a capitalization-weighted benchmark, macroscopic properties are clearly relevant since they affect both the absolute performance of the strategy and the relative performance with respect to the benchmark. For example, in [2, 46] it was shown empirically that changes in market diversity (Section 3.2), which measures the concentration of the market capitalization weights, correlate significantly with the relative return of actively managed large-cap portfolios. Macroscopic properties are also related to the overall stability of the market [22]. Viewing equity markets from the macroscopic perspective suggests many interesting questions which are not satisfactorily captured by conventional models of asset prices.
@@
Our study is motivated by stochastic portfolio theory (SPT) which was first pioneered by Fernholz [46, 50, 57]. This mathematical theory provides a fresh perspective on some macroscopic properties of large equity markets, especially those defined in terms of market capitalizations and/or returns. In particular, market diversity and intrinsic volatility (relative volatility among the stocks, see Section 4), appear to be stable over long periods and, at least under suitable conditions, can be exploited by carefully chosen portfolios to outperform a capitalization-weighted benchmark. While the size factor has played an important role in empirical finance (see e.g. [42]), SPT places great emphasis on how (relative) sizes simplify the modelling of the market, describe its stability, and allow us to analyze in depth the relative performance of a wide class of systematically rebalanced portfolios. Unfortunately, many results motivated by SPT remain largely inaccessible to non-experts due to their technical nature and possibly some restrictions imposed by the mathematical models (such as the interacting particle systems discussed in Section 3.1 and the focus on functionally generated portfolios). Also, the majority of papers in SPT focus on theoretical developments. Many important questions, such as the joint modelling and prediction of diversity and volatility, and the calibration and diagnostics of high dimensional market models, are largely open.
@@
Our paper, whose title is inspired by [26], has three main objectives. First and foremost, we present a self-contained and unified empirical study of some macroscopic properties of the US equity market from 1962 to 2023, based on the CRSP US Stock Database, for a broad audience including financial mathematicians, econometricians, financial economists and portfolio managers. We substantiate, and sometimes modify, the “folklore observations” in SPT, thereby suggesting a collection of statistical and financial problems that we believe are interesting for the aforementioned communities. Especially, we highlight the observation that the apparent stability of the capital distribution curve is closely related to the inflow (e.g. IPO) and outflow (e.g. delisting) of stocks. To facilitate replication of our empirical results and/or modification of the conventions and parameters adopted, we have made our codes publicly available. [^2] See https://github.com/stevenacampbell/Macroscopic-Properties-of-Equity-Markets. The CRSP Data, which is licensed, must be obtained separately. Second, we perform a systematic empirical backtesting of various specifications of the diversity-weighted portfolio [52] which is a representative large cap, rule-based strategy of special interest in portfolio management in general, and in SPT in particular. We show how their performance, relative to a capitalization-weighted benchmark, depend on the (realized) macroscopic quantities, extending the empirical studies in [100, 106]. Our backtesting engine, which implements the algorithm in [100] for long-only portfolios under proportional transaction costs, is also available on our repository and may be useful to other researchers. Third, as a by-product of our exposition, we offer discussions on recent developments motivated by SPT, extending beyond those covered in the 2009 survey [55], again with the broader communities in mind.
@@ 3.1. The capital distribution curve
The capitalization weights define a probability vector with values in the unit simplex
@@
Using the capitalization weights, we can visualize more accurately and vividly the evolution of the CRSP universe. In Figure 2, we show with each green dot (left panel) the time and size of a stock when it is newly listed, and with each red dot (right panel) the same information for a stock which is delisted for an extended period. We observe that both “births” (entrances) and “deaths” (exits) occur frequently across a wide range of market capitalizations except at the very top, yet their intensities appear to fluctuate over time. The weights at exit time are overall more spread out than the weights at entrance. An interesting statistical problem is to relate these fluctuations with the underlying economy such as business cycles.
@@
In Figure 3 we plot the full capital distribution curve, which is one of the most iconic objects in SPT (see e.g. [46, Chapter 5]), from 1962 to 2023. The evolving dimension of $\mathcal{A}_{t}$ is inconvenient for conventional statistical methods. To avoid this issue, one may restrict to the subuniverse $\mathcal{A}_{t}^{K}$ consisting of the largest $K$ stocks of $\mathcal{A}_{t}$; [^5] Exact ties occur extremely rarely, if at all, and can be resolved arbitrarily. this amounts to renormalizing $(\mu^{\mathcal{A}_{t}}_{(1)}(t),\ldots,\mu^{\mathcal{A}_{t}}_{(K)}(t))$. The capital distribution curves relative to $\mathcal{A}^{K}_{t}$ with $K=1000$ are visualized collectively in Figure 4 as a surface over the (log rank)-date plane. A glance at Figures 3 and 4 should convince the reader that the capital distribution curve is a rather delicate object to model faithfully - either directly or through a market model which specifies the joint dynamics of all stocks. Here, we observe that the left end of the curve is more volatile (in log-scale) than the right end, even though smaller stocks are generally more volatile in absolute terms (see [18, Section 5.3] and Section 5). Nevertheless, the overall shapes of the capital distribution curves appear to be quite stable. Thus we state our first stylized fact, somewhat loosely, as follows:
@@
The apparent long-term stability, which we distinguish from the concept of (strict or second order) stationarity in time series analysis, of the shape of the capital distribution curve is striking and calls for explanation. As noted in [36, pages 101–103], this stability is neglected by standard multivariate financial models. To take a simple example, suppose stocks followed a multivariate Black-Scholes model as in the classical Merton problem [87]. Then the capital distribution would eventually degenerate into a point mass by the law of large numbers. If the capital distribution followed approximately a power law (or Pareto distribution) with exponent $\alpha>0$, i.e., $\mu^{\mathcal{A}_{t}}_{(k)}\approx\frac{1}{Z_{t}}k^{-\alpha}$ where $Z_{t}$ is a normalization constant, the capital distribution curve would look approximately like a straight line with slope $-\alpha$. The special case $\alpha=1$ is frequently called Zipf’s law. While power laws are ubiquitous in many rank–size distributions in economics (see e.g. [58, 60, 89, 101] and the references therein) and complex systems [107], it does not hold for the full capital distribution curve but may be a reasonable first approximation. Here, we should mention that the CRSP universe does not contain private companies which are not listed on stock exchanges. We also remark that because the capital distribution curve is plotted in log-log scale, a visually small change of the curve, especially on its right end, correspond to large movements in the capitalization weights.
@@
Consider the problem of constructing stochastic models of market capitalizations which are capable of reproducing some features of the observed capital distribution curve (and ideally other behaviours of the market). A reasonably realistic model not only improves our understanding of the market but also serves as a scenario generator for optimizing and backtesting trading and risk management strategies. To enforce stability, a far-reaching idea, first proposed by Fernholz in [46, Section 5.5] based on empirical observations, is to let the dynamics of a stock depend on its rank. This modelling assumption motivated novel rank-based diffusions, and related processes such as volatility-stabilized processes and polynomial processes, for which a substantial probabilistic literature has accumulated in the last two decades (see e.g. [8, 25, 30, 32, 54, 58, 64, 65, 70, 79, 90, 103, 104] and their references). While it is beyond the scope of this paper to describe the variety of mathematical results obtained, we will explain the main ideas.
For theoretical tractability, one typically considers an equity market consisting of a fixed (finite or countably infinite) collection of stocks; in particular, IPOs and delisting events are neglected even though they are prominent as seen in Figure 1. Nevertheless, by considering the largest $K\leq n$ stocks of the system, one can obtain a simulated universe which evolves over time via rank switchings. As a basic example, we mention the generalized Atlas model, introduced in [8], under which the market capitalizations $X_{1}(t),\ldots,X_{n}(t)$ of $n$ stocks are modelled in continuous time as a Markov diffusion process satisfying the following system of stochastic differential equations:
@@
where $\mathsf{rank}_{i}(t)$ is the rank of stock $i$ at time $t$, $\gamma_{k}$ and $\sigma_{k}$ are suitable rank-based constants and $W_{1},\ldots,W_{n}$ are independent Brownian motions. In words, (3.3) says that the log market capitalization, $\log X_{i}$, of stock $i$ has drift $\gamma_{k}$ and volatility $\sigma_{k}$ when it has rank $k$ at time $t$. Focusing on rank-based behaviours limits the number of parameters. For example, the system (3.3) is specified by $2n$ parameters while an unrestricted covariance matrix requires $\frac{1}{2}n(n+1)$. In Section 5 we report some empirical evidence of rank dependence. Interacting particle systems such as (3.3) are not straightforward to analyze since the coefficients are discontinuous in the state variable. [^6] In (3.3) the coefficients are piecewise constant. Existence of a weak solution follows from the theory in [10]. Also, the dynamics of the ranked market weights $\mu_{(k)}(t)$ and the gaps $\log\mu_{(k)}(t)-\log\mu_{(k+1)}(t)$ involve delicate concepts such as local times, collisions, as well as, reflections over polyhedral boundaries. Under suitable structural conditions on the parameters, it can be shown that a process such as (3.3) is ergodic and has a unique stationary distribution, under which the (expectation of the) capital-distribution curve is qualitatively similar to the ones shown in Figure 3.777The subsystem consisting of the largest $K$ stock is called an open market in [74]. They also capture some other rank-based properties such as the higher volatility of smaller stocks (Section 5.1). Nevertheless, we observe that most, if not all, results in this literature rely on ergodicity in the limit $t\rightarrow\infty$ and the stationary distribution. Consequently, these models need not, and usually do not, capture (qualitatively or quantitatively) short and medium term behaviours of the observed equity market.
@@
The model in (3.3) is called a first order model since the coefficients depend only on the ranks. A second order or hybrid model, as in [53, 65, 70] involves also name-based coefficients and can incorporate idiosyncratic features. Calibration of ranked-based and volatility-stabilized market models to market data, as done in [46, 54, 64, 65, 69], is typically based on ad hoc methods analogous to moment matching and elementary estimates of growth rates and volatility coefficients (with smoothing); accuracy of the estimates and diagnostic checks are not addressed. A natural problem is to improve these models, their estimation and predictive power, possibly using tools in machine learning such as neural stochastic differential equations (see e.g. [82]). Moreover, market models such as (3.3) are exogenous; they do not explain why and how the stated dynamics emerge from interactions of the investors and the underlying economy. An ambitious goal is to come up with economically sound models of the financial market under which the observed macroscopic features emerge. Ideas from evolutionary finance [14, 41], mean-field games [21, 66, 72, 80] and complex systems [107] may be relevant in modelling the dynamics of firms and interactions among investors and/or their strategies.
@@ 3.2. Market diversity
By symmetry of the entropy we have ${\bf H}(\mu^{\mathcal{I}}(t))={\bf H}(\mu^{\mathcal{I}_{t}}_{\geq}(t)$). For a universe with a fixed size $|\mathcal{I}_{t}|$, ${\bf H}(\mu(t))$ is maximized (with value $\log|\mathcal{I}_{t}|$) when $\mu(t)$ is the uniform distribution, i.e., $\mu_{i}^{\mathcal{I}_{t}}(t)=\frac{1}{|\mathcal{I}_{t}|}$, and is minimized (with value 00) when some stock completely dominates the universe, i.e., $\mu_{i}^{\mathcal{I}_{t}}(t)=1$ for some $i$ and $\mu_{j}^{\mathcal{I}_{t}}(t)=0$ for $j\neq i$. Apart from the Shannon entropy, other diversity measures (see [46, Section 3.4]) may be used and produce qualitatively similar results; a specific example is the parameterized diversity measure ${\bf D}_{p}$ defined by
@@
where $p\in(0,1)$ is a tuning parameter. It is closely related to the Rényi entropy (see [114, Proposition 2]) and the diversity-weighted portfolio (A.7) whose performance will be backtested, under various specifications, in Section 6. In fact, under the classical setting in SPT, the change in $\log{\bf D}_{p}$ is a major component of the relative performance of the diversity-weighted portfolio with respect to the market portfolio, in the short to medium term (see (A.8)). More generally, in [2] (also see [51, 52]) it was shown that the diversity of the S&P 500 – which corresponds roughly to $\mathcal{A}_{t}^{500}$ – correlates significantly with the average relative performance of large-cap managers. Additional empirical results will be reported in Section 6.
@@
We show that the behaviours of market diversity depend crucially on the choice of the universe $\mathcal{I}_{t}$. On the other hand, that diversity is “mean reverting” is subject to doubt. In the strictest sense, mean reversion means diversity is well approximated by an autoregressive or Ornstein-Uhlenbeck process about a fixed equilibrium level. While our empirical results do not provide conclusive evidence of this claim, long-term probabilistic behaviours of diversity and other observables can be analyzed mathematically for rank-based systems which generalize (3.3); see [88].
@@
Observe that in 2023, even after the COVID-19 recession, the diversity of $\mathcal{A}_{t}$ dropped to the lowest level since the 1970s. The recent decrease in market diversity, largely attributable to seven mega-capitalization stocks, dubbed the “Magnificent Seven”, is affecting active managers’ stock selection behaviour. The active-share ratio, which measures how active portfolios deviate from the benchmark S&P 500, is at its lowest level since 2013 [110], indicating that active portfolio managers are choosing to mimic the benchmark more because missing out on any of the top-weighted stocks can lead to underperformance.
@@
    (Black series) $\mathcal{I}_{t}=\mathcal{A}^{K}_{t}$ is the largest $K$ stocks on day $t$ (renewed every day). Now diversity fluctuates in a much tighter interval compared to the case of full universe, but even with this restriction it is difficult to argue the existence of a fixed equilibrium value. Note that during the the period 1962–1972 the CRSP universe was artificially enlarged twice due to the addition of NYSE American and NASDAQ stocks. The diversity of $\mathcal{A}^{K}_{t}$ increased rather steadily in this period. We also observe that diversity appears to exhibit short to medium term trends or momentum. Also, several sudden drops in market diversity are associated with market crashes (such as the financial crisis in 2008 and COVID-19 in 2020).

### v2

@@ 1. Introduction
In this paper we study empirically a collection of macroscopic properties of equity markets which are highly relevant in the management of large equity portfolios but are not adequately addressed by the aforementioned literature. In a nutshell, a macroscopic quantity or property of an equity market is one which depends on all or a majority of the stocks in the market. The most basic example is the total market capitalization, or overall performance, of the market which is well represented by a capitalization-weighted market index such as the S&P 500. The index often serves as a benchmark for equity portfolios, both passive and active, and may be approximated by tradeable assets such as exchange-traded funds. Our terminology is motivated by statistical physics in which one considers macroscopic quantities such as volume, temperature and pressure. To take our crude physical analogy a bit further, suppose we think of an equity market as a galaxy and each stock as a star. Then the market index is a proxy of the total mass, but there are other interesting macroscopic quantities, such as the shape, spiral and rotation of the galaxy, which affect and are affected by the stars, and are much less understood. [^1] There are also other galaxies (markets) but they are beyond the scope of this paper. From the viewpoint of, say, the manager of an equity portfolio which aims to outperform a capitalization-weighted benchmark, macroscopic properties are clearly relevant since they affect both the absolute performance of the strategy and the relative performance with respect to the benchmark. For example, in [2, 47] it was shown empirically that changes in market diversity (Section 3.2), which measures the concentration of the market capitalization weights, correlate significantly with the relative return of actively managed large-cap portfolios. Macroscopic properties are also related to the overall stability of the market [23]. Viewing equity markets from the macroscopic perspective suggests many interesting questions which are not satisfactorily captured by conventional models of asset prices. The study of macroscopic features and their relationship with individual actions also arises in other areas in mathematical finance and applied probability, most notably in mean-field games [21, 66, 67, 85].
@@
Our study is motivated by stochastic portfolio theory (SPT) which was first pioneered by Fernholz [47, 51, 58]. This mathematical theory provides a fresh perspective on some macroscopic properties of large equity markets, especially those defined in terms of market capitalizations and/or returns. In particular, market diversity and intrinsic volatility (relative volatility among the stocks, see Section 4), appear to be stable over long periods and, at least under suitable conditions, can be exploited by carefully chosen portfolios to outperform a capitalization-weighted benchmark. While the size factor has played an important role in empirical finance (see e.g. [43]), SPT places great emphasis on how (relative) sizes simplify the modelling of the market, describe its stability, and allow us to analyze in depth the relative performance of a wide class of systematically rebalanced portfolios. Unfortunately, many results motivated by SPT remain largely inaccessible to non-experts due to their technical nature and possibly some restrictions imposed by the mathematical models (such as the interacting particle systems discussed in Section 3.1 and the focus on functionally generated portfolios). Also, the majority of papers in SPT focus on theoretical developments. Many important empirical questions, such as the joint modelling and prediction of diversity and volatility, and the calibration and diagnostics of high dimensional market models, are largely open.
Our paper, whose title is inspired by [27], has three main objectives. First and foremost, we present a self-contained and unified empirical study of some macroscopic properties of the US equity market from 1962 to 2023, based on the CRSP US Stock Database, for a broad audience including financial mathematicians, econometricians, financial economists and portfolio managers. We substantiate, and sometimes modify, the “folklore observations” in SPT, thereby suggesting a collection of statistical and financial problems that we believe are of interest to the aforementioned communities. We summarize our main empirical findings which will be made more precise in the body of the paper:
  • 1.
    Capital distribution curve is stable: The dependence of market capitalization weight on relative rank (by size) does not follow Zipf’s law but has been stable over decades.
  • 2.
    Market diversity is maintained through entrances and exits of stocks. The capital distribution of a fixed collection of stocks tends to become more concentrated over time.
  • 3.
    Cross-sectional volatility clusters in time and correlates with change in market diversity: The relative volatility among stocks exhibits ARCH effects and tends to be high when market concentration changes quickly.
  • 4.
    Smaller stocks are more volatile and tend to become (relatively) smaller. Smaller stocks have systematically larger absolute volatilities, cross ranks more often, and has a higher tendency to be overtaken by smaller or new stocks.
@@
To facilitate replication of our empirical results and/or modification of the conventions and parameters adopted, we have made our codes publicly available. [^2] See https://github.com/stevenacampbell/Macroscopic-Properties-of-Equity-Markets. The CRSP Data, which is licensed, must be obtained separately. Second, we perform a systematic empirical backtesting of various specifications of the diversity-weighted portfolio [53] which is a representative large cap, rule-based strategy of special interest in portfolio management in general, and in SPT in particular. We show how their performance, relative to a capitalization-weighted benchmark, depend on the (realized) macroscopic quantities, extending the empirical studies in [105, 111]. Our backtesting engine, which implements the algorithm in [105] for long-only portfolios under proportional transaction costs, is also available on our repository and may be useful to other researchers and practitioners. Third, as a by-product of our exposition, we offer discussions on recent developments motivated by SPT, extending beyond those covered in the 2009 survey [56], again with the broader communities in mind.
@@ 3.1. The capital distribution curve
The capitalization weights define a probability vector with values in the (closed) unit simplex
@@
Using the capitalization weights, we can visualize more accurately and vividly the evolution of the CRSP universe. In Figure 2, we show with each green dot (left panel) the time and size of a stock when it is newly listed, and with each red dot (right panel) the same information for a stock which is delisted forever or for an extended period. We observe that both “births” (entrances) and “deaths” (exits) occur frequently across a wide range of market capitalizations, yet their intensities appear to fluctuate over time. The weights at exit time are overall more spread out than the weights at entrance. An interesting statistical problem is to relate these fluctuations with the underlying economy such as business cycles.
@@
In Figure 3 we plot the full capital distribution curve, which is one of the most iconic objects in SPT (see e.g. [47, Chapter 5]), from 1962 to 2023. The evolving dimension of $\mathcal{A}_{t}$ is inconvenient for conventional statistical methods. To avoid this issue, one may restrict to the subuniverse $\mathcal{A}_{t}^{K}$ consisting of the largest $K$ stocks of $\mathcal{A}_{t}$; [^5] Exact ties occur extremely rarely, if at all, and can be resolved arbitrarily. this amounts to renormalizing $(\mu^{\mathcal{A}_{t}}_{(1)}(t),\ldots,\mu^{\mathcal{A}_{t}}_{(K)}(t))$. The capital distribution curves relative to $\mathcal{A}^{K}_{t}$ with $K=1000$ are visualized collectively in Figure 4 as a surface over the (log rank)-date plane. A glance at Figures 3 and 4 should convince the reader that the capital distribution curve is a rather delicate object to model faithfully - either directly or through a market model which specifies the joint dynamics of all stocks. Here, we observe that the left end of the curve is more volatile (in log-scale) than the right end, even though smaller stocks are generally more volatile in absolute terms (see [18, Section 5.3] and Section 5). Nevertheless, the overall shapes of the capital distribution curves appear to be quite stable. Thus we state our first stylized fact, somewhat loosely, as follows:
@@
The apparent long-term stability, which we distinguish from the concept of (strict or second order) stationarity in time series analysis, of the shape of the capital distribution curve is striking and calls for explanation. As noted in [37, pages 101–103], this stability is neglected by standard multivariate financial models. To take a simple example, suppose stocks followed a multivariate Black-Scholes model as in the classical Merton problem [92]. Then the capital distribution would eventually degenerate into a point mass by the law of large numbers. If the capital distribution followed approximately a power law (or Pareto distribution) with exponent $\alpha>0$, i.e., $\mu^{\mathcal{A}_{t}}_{(k)}\approx\frac{1}{Z_{t}}k^{-\alpha}$ where $Z_{t}$ is a normalization constant, the capital distribution curve would look approximately like a straight line with slope $-\alpha$. The special case $\alpha=1$ is frequently called Zipf’s law. While power laws are ubiquitous in many rank–size distributions in economics (see e.g. [59, 62, 94, 106] and the references therein) and complex systems [112], it does not hold for the full capital distribution curve but may be a reasonable first approximation. Here, we should mention that the CRSP universe does not contain private companies which are not listed on stock exchanges. We also remark that because the capital distribution curve is plotted in log-log scale, a visually small change of the curve, especially on its right end, correspond to large movements in the capitalization weights.
@@
Consider the problem of constructing stochastic models of market capitalizations which are capable of reproducing some features of the observed capital distribution curve (and ideally other behaviours of the market). A reasonably realistic model not only improves our understanding of the market but also serves as a scenario generator for optimizing and backtesting trading and risk management strategies. To enforce stability, a far-reaching idea, first proposed by Fernholz in [47, Section 5.5] based on empirical observations, is to let the dynamics of a stock depend on its rank. This modelling assumption motivated novel rank-based diffusions, and related processes such as volatility-stabilized processes and polynomial processes, for which a substantial probabilistic literature has accumulated in the last two decades (see e.g. [8, 26, 31, 33, 55, 59, 68, 69, 74, 84, 95, 108, 109] and their references). While it is beyond the scope of this paper to describe the variety of mathematical results obtained, we will explain the main ideas.
For theoretical tractability, one typically considers an equity market consisting of a fixed (finite or countably infinite) collection of stocks; in particular, IPOs and delisting events are neglected even though they are prominent as seen in Figure 1.666A notable exception is [80] which studied models with splits and mergers. Nevertheless, by considering the largest $K\leq n$ stocks of the system, one can obtain a simulated universe which evolves over time via rank switchings. [^7] The subsystem consisting of the largest $K$ stock is called an open market in [78]. As a basic example, we mention the generalized Atlas model, introduced in [8], under which the market capitalizations $X_{1}(t),\ldots,X_{n}(t)$ of $n$ stocks are modelled in continuous time as a Markov diffusion process satisfying the following system of stochastic differential equations:
@@
where $\mathsf{rank}_{i}(t)$ is the rank of stock $i$ at time $t$, $\gamma_{k}$ and $\sigma_{k}$ are suitable rank-based constants and $W_{1},\ldots,W_{n}$ are independent Brownian motions. In words, (3.3) says that the log market capitalization, $\log X_{i}$, of stock $i$ has drift $\gamma_{k}$ and volatility $\sigma_{k}$ when it has rank $k$ at time $t$. Focusing on rank-based behaviours limits the number of parameters. For example, the system (3.3) is specified by $2n$ parameters while an unrestricted covariance matrix requires $\frac{1}{2}n(n+1)$. In Section 5 we report some empirical evidence of rank dependence. Interacting particle systems such as (3.3) are not straightforward to analyze since the coefficients are discontinuous in the state variable. [^8] In (3.3) the coefficients are piecewise constant. Existence of a weak solution follows from the theory in [10]. Also, the dynamics of the ranked market weights $\mu_{(k)}(t)$ and the gaps $\log\mu_{(k)}(t)-\log\mu_{(k+1)}(t)$ involve delicate concepts such as local times, collisions, as well as, reflections over polyhedral boundaries. Under suitable structural conditions on the parameters, it can be shown that a process such as (3.3) is ergodic and has a unique stationary distribution, under which the (expectation of the) capital-distribution curve is qualitatively similar to the ones shown in Figure 3. They also capture some other rank-based properties such as the higher volatility of smaller stocks (Section 5.1). Nevertheless, we observe that most, if not all, results in this literature rely on ergodicity in the limit $t\rightarrow\infty$ and the stationary distribution. Consequently, these models need not, and usually do not, capture (qualitatively or quantitatively) short and medium term behaviours of the observed equity market.
@@
The model in (3.3) is called a first order model since the coefficients depend only on the ranks. A second order or hybrid model, as in [54, 69, 74] involves also name-based coefficients and can incorporate idiosyncratic features. Calibration of ranked-based and volatility-stabilized market models to market data, as done in [47, 55, 68, 69, 73], is typically based on ad hoc methods analogous to moment matching and elementary estimates of growth rates and volatility coefficients (with smoothing); accuracy of the estimates and diagnostic checks are not addressed. A natural problem is to improve these models, their calibration and predictive power, possibly using tools in machine learning such as neural stochastic differential equations (see e.g. [87]). Moreover, market models such as (3.3) are exogenous; they do not explain why and how the stated dynamics emerge from interactions of the investors and the underlying economy. An ambitious goal is to come up with economically sound models of the financial market under which the observed macroscopic features emerge. Ideas from evolutionary finance [14, 42], mean-field games [22, 70, 76, 85] and complex systems [112] may be relevant in modelling the dynamics of firms and interactions among investors and/or their strategies.
@@ 3.2. Market diversity
By symmetry of the entropy we have ${\bf H}(\mu^{\mathcal{I}}(t))={\bf H}(\mu^{\mathcal{I}_{t}}_{\geq}(t)$). For a universe with a fixed size $|\mathcal{I}_{t}|$, ${\bf H}(\mu(t))$ is maximized (with value $\log|\mathcal{I}_{t}|$) when $\mu(t)$ is the uniform distribution, i.e., $\mu_{i}^{\mathcal{I}_{t}}(t)=\frac{1}{|\mathcal{I}_{t}|}$, and is minimized (with value 00) when some stock completely dominates the universe, i.e., $\mu_{i}^{\mathcal{I}_{t}}(t)=1$ for some $i$ and $\mu_{j}^{\mathcal{I}_{t}}(t)=0$ for $j\neq i$. Apart from the Shannon entropy, other diversity measures (see [47, Section 3.4]) may be used and produce qualitatively similar results; a specific example is the parameterized diversity measure ${\bf D}_{p}$ defined by
@@
where $p\in(0,1)$ is a tuning parameter. It is closely related to the Rényi entropy (see [119, Proposition 2]) and the diversity-weighted portfolio (A.7) whose performance will be backtested, under various specifications, in Section 6. In fact, under the classical setting in SPT, the change in $\log{\bf D}_{p}$ is a major component of the relative performance of the diversity-weighted portfolio with respect to the market portfolio, in the short to medium term (see (A.8)). More generally, in [2] (also see [52, 53]) it was shown that the diversity of the S&P 500 – which corresponds roughly to $\mathcal{A}_{t}^{500}$ – correlates significantly with the average relative performance of large-cap managers. Additional empirical results will be reported in Section 6.
@@
In the following we show that the behaviours of market diversity depend crucially on the choice of the universe $\mathcal{I}_{t}$. On the other hand, that diversity is “mean reverting” is subject to doubt. In the strictest sense, mean reversion means diversity is well approximated by an autoregressive or Ornstein-Uhlenbeck process about a fixed (or possibly varying) equilibrium level. While our empirical results do not provide conclusive evidence of this claim, long-term probabilistic behaviours of diversity and other observables can be analyzed mathematically for rank-based systems which generalize (3.3); see [93].
@@
Observe that in 2023, even after the COVID-19 recession, the diversity of $\mathcal{A}_{t}$ dropped to the lowest level since the 1970s. The recent decrease in market diversity, largely attributable to seven mega-capitalization stocks, dubbed the “Magnificent Seven”, is affecting active managers’ stock selection behaviour. The active-share ratio, which measures how active portfolios deviate from the benchmark S&P 500, is at its lowest level since 2013 [115], indicating that active portfolio managers are choosing to mimic the benchmark more because missing out on any of the top-weighted stocks can lead to underperformance. If this trend continues, the (left end of the) capital distribution curve will look slightly steeper than before.
@@
    (Black series) $\mathcal{I}_{t}=\mathcal{A}^{K}_{t}$ is the largest $K$ stocks on day $t$ renewed every day. Now diversity fluctuates in a much tighter interval compared to the case of full universe, but even with this restriction it is difficult to argue the existence of a fixed equilibrium value. Note that during the the period 1962–1972 the CRSP universe was artificially enlarged twice due to the addition of NYSE American and NASDAQ stocks. The diversity of $\mathcal{A}^{K}_{t}$ increased rather steadily in this period. We also observe that diversity appears to exhibit short to medium term trends or momentum. Also, several sudden drops in market diversity are associated with market crashes such as the financial crisis in 2008 and COVID-19 in 2020.